1.What is the final concentration of a solution of KCl that is prepared by adding 46mL of a 2.98M KCl solution to a flask for a final volume of 245 mL?

2.What volume( in mL) of a 1.89M solution of NaF is needed to prepare 406mL of a 0.35M solution of NaF?

3.Calculate the mass of Al metal needed to react with 87 mL of a 3.18M solution of HNO3. Al(s) + HNO3(aq) --> Al(NO3)3 + H2(g) ( unbalanced)

4.Calculate the mass of product formed when 135 mL of a 0.45M Fe(NO3)3 reacts with 75mL of a 0.75M solution of Ba(OH)2. Fe(NO3)3(aq) + Ba(OH)2 --> Ba(NO3)2(aq) + Fe(OH)3(s) (unbalanced)

5.Automobile batteries use 3.0 M H2SO4 as an electrolyte. How much 1.20 M NaOH will be needed to neutralize 225 mL of battery acid? H2SO4(aq) + 2NaOH(aq) → 2H2O(l) + Na2SO4(aq)
Answer #1

1)

use dilution formula

M1*V1 = M2*V2

1---> is for stock solution

2---> is for diluted solution

Given:

M1 = 2.98 M

V1 = 46 mL

V2 = 245 mL

use:

M1*V1 = M2*V2

M2 = (M1*V1)/V2

M2 = (2.98*46)/245

M2 = 0.560 M

Answer: 0.560 M

Only 1 question at a time please